# Stability of heat kernel bounds under pointed Gromov-Hausdorff convergence

### v1

@@ 8 Some examples
    We denote $d_{\mathbb{R}^{2}}$ be a metric on $\mathbb{R}^{2}$ (and on its subsets) defined by $d_{\mathbb{R}^{2}}((x_{1},y_{1}),(x_{2},y_{2}))=\max(|x_{1}-x_{2}|,|y_{1}-y_{2}|)$. Let $d_{0}$ (resp. $d_{\infty}$) be the intrinsic distance of $X_{0}$ (resp. $X_{\infty}$) with respect to $d_{\mathbb{R}^{2}}$, i.e. $d_{0}(x,y)$ (resp. $d_{\infty}(x,y)$) is the length of the shortest path in $X$ (resp. $X_{\infty}$) connecting $x$ and $y$. It is known that $(X_{0},d_{0})$ and $(X_{\infty},d_{\infty})$ are complete separable geodesic metric spaces, Moreover, $d_{0}$ and ${d_{\mathbb{R}^{2}}}|_{X_{0}}$, $d_{\infty}$ and ${d_{\mathbb{R}^{2}}}|_{X_{\infty}}$ are bi-Lipschitz equivalent [Bar13].
@@
    $\mathcal{E}_{0}(u,v):=\int_{\mathcal{P}}\nabla u\nabla v\mathop{}\!\mathrm{d}m_{0},\ u,v\in\mathcal{F}_{0}:=W^{1,2}(X_{0})$
@@
Barlow in [Bar04] shows that, the existence of an infinite locally finite weighted graph that satisfying $\mathrm{V}_{\alpha}$, $\mathrm{EHI}$ and $\mathrm{E}_{\beta}$, which is equivalent to the random walk satisfying the following type heat kernel estimates: for some $C_{1},c_{2}\geq 1$ and all $n\geq 1$,

### v2

@@ 8 Some examples
    We denote $d_{\mathbb{R}^{2}}$ be a metric on $\mathbb{R}^{2}$ (and on its subsets) defined by $d_{\mathbb{R}^{2}}((x_{1},y_{1}),(x_{2},y_{2}))=|x_{1}-x_{2}|+|y_{1}-y_{2}|\asymp\max(|x_{1}-x_{2}|,|y_{1}-y_{2}|)$. Let $d_{0}$ (resp. $d_{\infty}$) be the intrinsic distance of $X_{0}$ (resp. $X_{\infty}$) with respect to $d_{\mathbb{R}^{2}}$, i.e. $d_{0}(x,y)$ (resp. $d_{\infty}(x,y)$) is the length of the shortest path in $X$ (resp. $X_{\infty}$) connecting $x$ and $y$. It is known that $(X_{0},d_{0})$ and $(X_{\infty},d_{\infty})$ are complete separable geodesic metric spaces. Moreover, $d_{0}$ and ${d_{\mathbb{R}^{2}}}|_{X_{0}}$, $d_{\infty}$ and ${d_{\mathbb{R}^{2}}}|_{X_{\infty}}$ are bi-Lipschitz equivalent [Bar13].
@@
    $\mathcal{E}_{0}(u,v):=\int_{\mathcal{P}}\nabla u\cdot\nabla v\mathop{}\!\mathrm{d}m_{0},\ u,v\in\mathcal{F}_{0}:=W^{1,2}(X_{0})$
@@
Barlow in [Bar04] shows that, the existence of an infinite locally finite weighted graph that satisfying $\mathrm{V}_{\alpha}$, $\mathrm{EHI}$ and $\mathrm{E}_{\beta}$ (see [Bar04] for definitions), which is equivalent to the random walk satisfying the following type heat kernel estimates: for some $C_{1},c_{2}\geq 1$ and all $n\geq 1$,